Find out the frequency and wave number if energy radiation wavelength of 200 angstrom

Find  out the frequency and wave number if energy radiation wavelength of 200 angstrom 

Grade:11

2 Answers

birbal kushwaha
5 years ago
Frequency,  f =c/wavelenth= 3*10^8/200*10^-10= 1.5*10^16 Hrz
Wave number : it tells that how many complete wave originated in one second
Wave number = 1/f
akash dutta
5 years ago
wave number=1/wavelength=1/200 A=1/200×10^-10=10^10×1/200=10^8×1/2
=0.5×10^8
v=c/w
 =3×10^8/0.5×10^8=6Hz
TThanks !!!